# Rate-optimal and computationally efficient nonparametric estimation on the circle and the sphere

### v1

@@ 5.1 The uniform distribution
9:   $a_{ij}\leftarrow Z_{j}/\sqrt{Z_{1}+\dots+Z_{d+1}}$
@@ 5.2 The von Mises-Fisher’s distribution
$f(x)=c_{d,\kappa}\exp\big(\kappa\langle x,\mu\rangle\big),\quad x\in\mathbb{S}^{d},$ (5.3)
@@
We can see that all values of $s$ gave very accurate estimations of the probabilities over each region. The plots and MISE both indicate that $s=2$ gives the best fit out of the values we considered. See Figure 4 for the MISE as $s$ ranges over $\{0.1,0.2,\dots,10\}$ where we see an approximate local minimum at $s=2.3$. We will see in the next examples that smaller values of $s$ will be more appropriate for densities with more local behaviour, and we will later use this conclusion in Section 6.3.
@@
Region $s=0.5$ $s=1$ $s=2$ Frequency True Prob.
@@
Unlike the last two examples, both the MISE and plot show that $s=2$ is too large to provide a good fit in this case. This is expected as the density has more local behaviour since it is a mixture density and each component vMF density has a higher shape parameter than before. Indeed, the $s=1$ plot and MISE indicate a better performance than $s=2$, with $s=0.5$ appearing as the best of the three. Moreover, the $s=2$ probabilities are much less accurate than the $s=0.5$ estimations, especially over smaller areas. This is attributed to the oversmoothing we see in this $s=2$ case.
@@
We summarize the findings of the Section. We have demonstrated the efficacy of our finite KDEs and probability estimation methods in several examples. The outcomes from these simulations were all in line with the theoretical results presented earlier. Moreover, as we noticed, small differences in the values of the selected $s$ give very similar results.
Very large values of $s$, is better to be avoided in a data analysis, because tend to oversmooth.
Very local phenomena are better expressed with values of $s<1$.
With theoretical and empirical evidence for the accuracy of these estimators in hand, we now turn to several case studies with real-world datasets in order to demonstrate how these methods can be used in several research fields.
@@ 6.1 Spatially Periodic Data: Honeybee Orientations
In this dataset, bees are labelled as 0, if their orientation is clear, or $1$, if their body is partially obscured within the hive. For consistency in our study, we only look at the $n=691$ honeybee orientations of class 0, which are given in degrees. Converting these to radians makes these observations suitable for a KDE as in Theorem 3.1 and Remark 3.1. We selected $s=1$, and therefore $h=n^{-1/(2s+1)}\approx 0.113$, $N_{s}=14$ and $r=5$. The data and corresponding KDE are plotted in Figure 8. In this frame, the bees exhibit a clear preference for facing approximately towards the angle 0.09-0.09 radians with the least favoured p
@@ 6.2 Temporally Periodic Data: Annual Precipitation Patterns
More generally, if a random variable $T$ is periodic on $(a,b]$ with density $f_{T}:(a,b]\to[0,\infty)$ satisfying $f_{T}(a)=f_{T}(b)$, then independent observations $T_{1},\dots,T_{n}\in(a,b]$ may be mapped to circular data on $\mathbb{S}^{1}$ by
$(a,b]$ $\to(-\pi,\pi],$
@@
With this converted data $\theta_{1},\dots,\theta_{1455}$ in hand, we construct a KDE as in (3.3), taking $s=2$, $h\approx 0.233$, $N_{s}=6$ as in Theorem 3.1, and $r=6$. The data, alongside such a KDE is plotted in Figure 9 with the first day of each month indicated on the plot. We can see that the wet spring season is clearly highlighted by the KDE as well as the drier summers with less rain.
@@ 6.3 Global Data: Earthquake Locations
With the above conversion in hand, we have observations $X_{1},\dots,X_{1603}\in\mathbb{S}^{2}$. We select $s=0.05$ and construct a KDE $\widehat{f}_{s}$, as in Theorem 3.1 and Remark 3.1, taking $h=n^{-1/(2s+2)}\approx 0.0295$, $N_{s}=92$ and $r=6$. The corresponding KDE as well as the data is plotted in Figure 10.
@@
Country Min. Lat. Max. Lat. Min. Lon. Max. Lon. Frequency KDE Prob.
@@ 6.4 Astronomical Data: Star Directions
Region Min. Lat Max. Lat. Min. Lon. Max. Lon. Frequency KDE Prob.
@@ 7 Conclusion
In our study we provided nonparametric density and probability estimators on the circle $\mathbb{S}^{1}$ and the sphere $\mathbb{S}^{2}$ that are both rate-optimal and computationally efficient. The key idea is to replace the usual infinite expansions underlying kernel density estimators with their finite-order counterparts, determining the order index in accordance with the smoothness of the density.
@@
There are several natural directions for future research. One is the extension of the finite-order methodology to higher-dimensional spheres $\mathbb{S}^{d}$, $d>2$, or other metric spaces associated with Laplacians with discrete spectrum as in [10]. Another very natural direction is the study of adaptive adaptive estimators that do not rely on prior smoothness assumptions, something that is in our future agenda. Finally, Bayesian nonparametric models for directional data may provide a promising route for future research.
@@ Acknowledgements
The first-named author would like to thank Prof. Morgan Fraser for his helpful comments and suggestions on the astronomical application discussed in Section 6.4. The second-named author would like to thank Prof. Dr. Katarzyna Bozek an author of [8] and [9], whose dataset and expertise enabled the discussion in zoology in Section 6.1.
@@ 8 Appendix
$\text{MSE}(\widehat{f}_{s})(x)$ $\leq 2\left(\mathbb{E}\Big[\big(f(x)-\widehat{f}_{n,h}(x)\big)^{2}\Big]+\mathbb{E}\Big[\big(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\big)^{2}\Big]\right)$ (8.1)
$=2\left(\text{MSE}(\widehat{f}_{n,h})(x)+\mathbb{E}\Big[\big(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\big)^{2}\Big]\right).$
@@
$\mathbb{E}\left[\left(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\right)^{2}\right]\leq n^{-2s/(2s+1)}.$
@@ 8.2 Proofs of the results in probability estimation
Then by the Cauchy-Schwarz inequality for $\mathcal{L}^{2}(\mathbb{S}^{d})$ and Fubini-Tonelli Theorem we have that
$\mathbb{E}_{f}$ $\big(\widehat{\mathbb{P}}(A)-\mathbb{P}(A)\big)^{2}\leq\mathbb{E}_{f}\Big(\int_{A}\big|\widehat{f}_{s}(x)-f(x)\big|d\sigma_{\mathbb{S}^{d}}(x)\Big)^{2}$
@@
We are concerned with evaluating $\int_{A[\vartheta_{1},\vartheta_{2}]}\widehat{f}(x)d\sigma_{\mathbb{S}^{1}}(x)=\int_{\vartheta_{1}}^{\vartheta_{2}}\widehat{f}_{s}(\theta)d\theta$ where $\widehat{f}_{s}$ is as in (3.3). By simple integration we confirm the statement:
$\widehat{\mathbb{P}}$ $(A[\vartheta_{1},\vartheta_{2}])=\int_{\vartheta_{1}}^{\vartheta_{2}}\frac{1}{2\pi n}\sum_{j=1}^{n}\left(1+2\sum_{\ell=1}^{N_{s}}g(h\ell)\cos(\ell(\theta-\theta_{j}))\right)d\theta$
@@
Therefore for the integral $\int_{\gamma_{1}}^{\gamma_{2}}P_{\ell}^{m}(u)du$, it suffices to compute $\int_{\gamma_{1}}^{\gamma_{2}}(1+u)^{m/2}(1-u)^{k-m/2}du$. We consider the more general integral
@@
$\text{Re}\big(Y^{m}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{m}(\theta_{j},\phi_{j})}\big)$ $=N_{\ell,m}^{2}P_{\ell}^{m}(\cos\theta)P^{m}_{\ell}(\cos\theta_{j})\text{Re}\Big(e^{im(\phi-\phi_{j})}\Big)$

### v2

@@ 5.1 The uniform distribution
9:   $a_{ij}\leftarrow Z_{j}/\sqrt{Z_{1}^{2}+\dots+Z_{d+1}^{2}}$
@@ 5.2 The von Mises-Fisher’s distribution
$f(x)=c_{d,\kappa}\exp\big(\kappa\langle x,\mu\rangle\big)=c_{d,\kappa}e^{2\kappa}\exp\big(-\frac{\kappa}{2}\|x-\mu\|^{2}\big),\quad x\in\mathbb{S}^{d},$ (5.3)
@@
We can see that all values of $s$ gave very accurate estimations of the probabilities over each region. The plots and MISE both indicate that $s=2$ gives the best fit out of the values we considered. See Figure 4 for the MISE as $s$ ranges over $\{0.1,0.2,\dots,10\}$ where we see an approximate local minimum at $s=2.3$.
@@
Scenario 4 (vMF on $\mathbb{S}^{1}$).
We simulate $n=1000$ data from the vMF distribution on the circle with $\kappa=10$ and $\mu=(0,1)$. The findings are summarized in Figure 5 and Table 4.
Figure 5: (Left to right) The plot of a vMF density over $\mathbb{S}^{1}$ with $\kappa=10$, $\mu=(0,1)$; a sample of $n=1000$ points from such a distribution; KDEs constructed using said data for $s=0.5$, $s=1$ and $s=2$, respectively.
Region $s=0.5$ $s=1$ $s=2$ Frequency True Prob.
$N_{s}=85$ $N_{s}=17$ $N_{s}=6$
$0,\pi]$ 0.999997 0.999944 0.999887 1 0.999989
$\pi/3,2\pi/3]$ 0.873325 0.874004 0.817978 0.874 0.893612
$2\pi/5,3\pi/5]$ 0.656978 0.649173 0.554804 0.662 0.671049
$5\pi/11,6\pi/11]$ 0.327823 0.328058 0.267136 0.329 0.343864
MISE 0.00762546 0.00214386 0.0445443 - -
Table 4: Estimated probabilities from a vMF density with $\kappa=10$ and $\mu=(0,1)$ on $\mathbb{S}^{1}$ for $n=1000$ points and $s=0.5$, $s=1$, $s=2$.
5.3 Mixture distributions
@@
Region $s=0.5$ $s=1$ $s=2$ Freq. True Prob.
@@
Unlike Scenarios 1 and 3, both the MISE and plot show that $s=2$ is too large to provide a good fit in this case. This is expected as the density has more local behaviour since each mixture component vMF density has a very high shape parameter. Indeed, the $s=1$ plot and MISE indicate a better performance than $s=2$, with $s=0.5$ appearing as the best of the three. Moreover, the $s=2$ probabilities are much less accurate than the $s=0.5$ estimations, especially over smaller areas.
Scenario 6 (vMF mixture on $\mathbb{S}^{1}$).
@@
Remark 5.1.
We summarize our conclusions as follows:
(a) Near-uniform settings, such as Scenarios 1 and 2, as well as Scenario 3—where the vMF peak for $\kappa=1$ is very low—are well estimated using larger values of $s$ (like $s\approx 2$), and equivalently a larger bandwidth $h$.
(b) Moderate values of $s$ (such as $s\approx 1$), and hence of $h$, are preferable when the underlying distribution departs from uniformity and exhibits more pronounced peaks along with flatter regions.
(c) The only case where we observed a very small value of $s=0.5$ was Scenario 5, which involves two high peaks ($12$ and $10$) separated by a distance of $\pi/2$, leading to localized concentration in distinct regions with extended flat areas between them.
5.4 Comparison with vMF estimators
Density estimation on the sphere dates back four decades. In particular, [26] has attracted significant attention and represents one of the most widely used methods for handling spherical data. The kernel density estimators proposed in [26, Section 5] have the form
$\widetilde{f}_{\kappa}(x):=\frac{c_{d,\kappa}}{n}\sum_{i=1}^{n}\exp\big(\kappa\langle x,X_{i}\rangle\big),\quad\kappa>0,$
where $\kappa>0$ is a smoothing parameter playing the role of the reciprocal of a bandwidth, and $c_{d,\kappa}$ is as in (5.3). In simple terms, the estimator $\widetilde{f}_{\kappa}$ averages vMF densities with common concentration parameter $\kappa$, centred at each observation $X_{i}$, and is often referred to as the vMF kernel density estimator (vMF-KDE).
In [26, Section 5], the parameter $\kappa$ is selected via cross-validation methods by minimizing the squared-error loss and the Kullback–Leibler loss, respectively. In this section, we compare our six simulation scenarios with the vMF-KDEs obtained via cross-validation methods in [26]. Each of the two loss functions yields an estimated value of the parameter $\kappa$, which we denote by $\hat{\kappa}_{2}$ and $\hat{\kappa}_{\mathrm{KL}}$, respectively, and gives rise to the corresponding vMF-KDEs, denoted by $\widetilde{f}_{\hat{\kappa}_{2}}$ and $\widetilde{f}_{\hat{\kappa}_{\mathrm{KL}}}$.
The estimated $\mathrm{MISE}$ for each estimator across all scenarios is reported in Table 7. As shown, the vMF-KDEs introduced in [26] achieve smaller $\mathrm{MISE}$ values than the finite-order estimators of Theorem 3.1 in Scenarios 1 and 2, corresponding to the uniform distribution, when the target density is constant.
In contrast, the finite-order estimators established here attain smaller $\mathrm{MISE}$ values in Scenarios 3–6, where the underlying densities (vMF distributions and their mixtures) exhibit one or more modes. This highlights the ability of our estimators to accurately capture local features of the density, including sharper extrema and appropriate decay in low-density regions; see also Figure 8.
Scenario $\widehat{f}_{s}$ $\widetilde{f}_{\hat{\kappa}_{2}}$ $\widetilde{f}_{\hat{\kappa}_{\mathrm{KL}}}$
[…]
Table 7: The estimated MISE of the KDE $\widehat{f}_{s}$, alongside the vMF-KDEs $\widetilde{f}_{\hat{\kappa}_{2}}$ and $\widetilde{f}_{\hat{\kappa}_{\mathrm{KL}}}$ for the value of $s>0$ that minimised MISE or the mean of the minimisers $\hat{\kappa}$ of each cross-validation function.
Figure 8: A closer view of the finite-order KDE in Scenario 5 in comparison to vMF-KDEs at the peak of the density (left) and the tail (right), plotted in terms of polar coordinates. The estimator $\widehat{f}_{s}$ captures both the peak and the tail.
5.5 Comparison with adaptive density estimators
Adaptive estimation is among the most important developments in nonparametric statistics and has been studied extensively over several decades across a variety of problems.
Adaptive density estimation on the sphere has been developed in [1, 2] and further used in [19], through wavelet estimators (note that wavelets are often referred to as needlets on compact manifolds), with coefficients selected via hard thresholding. This powerful methodology was introduced in [17] and later disseminated to a broader audience in the monograph [27]. In such adaptive estimators, there is no need to assume a given level $s$ of regularity, rather than a supremum value of the density $\|f\|_{\infty}$ and a parameter related with the threshold (parameter $k_{0}$ in [1]).
In [1, Section 6], adaptive wavelet estimators are numerically implemented for estimating a vMF mixture on $\mathbb{S}^{2}$, and compared with vMF-KDEs, using the (observed) $\mathcal{L}^{\infty}$-distance between the estimators and the target density.
In this Section we compare our finite-order estimators with the adaptive wavelet estimators in the same scenario as in [1]: consider a mixture of $m=2$ vMF densities of the form (5.4), with weights $W=(0.7,0.3)$, shape parameters $K=(1.2,8)$, and means $M=\big((0,1,0),\big(-\frac{\sqrt{2}}{2},-\frac{\sqrt{2}}{2},0\big)\big)$ and simulate $n=8000$ points from its density $f$.
We estimate the $\mathcal{L}^{\infty}$-distance of our finite-order estimators, with $s=1$, following our Remark 5.1, as the peaks are just moderately high. We refer further to the Figure 9, where one can see the estimated MISE for several values of $s$. The minimum estimated MISE came in the case of $s=1.25$ with comparable performance for $s=1$ and $s=1.5$.
Figure 9: Boxplots for the estimated MISE of the finite-order KDE with $s\in\{0.5,0.75,1,1.25,1.5,1.75,2\}$, in the scenario as in [1]. The estimated MISEs are $0.00324$, $0.00133$, 0.00070, 0.00050, 0.00076, 0.00166, 0.00326, respectively.
More precisely, we evaluate the error of the estimator $|f(x)-\widehat{f}_{s}(x)|$ on a $0.01\times 0.01$ resolution grid over $0,\pi]\times(-\pi,\pi]$, in spherical coordinates, giving approximately $2\times 10^{5}$ evaluations. For a given simulation, we record the maximum of the error $|f(x)-\widehat{f}_{s}(x)|$ over this grid. This is performed $30$ times and the $\mathcal{L}^{\infty}$-distance is estimated as the mean of these observed maximum errors. The results from this study, as well as the estimated $\mathcal{L}^{\infty}$-distance of the wavelet estimators for $k_{0}\in\{0.25,0.45,0.7\}$ are given in Table 8.
Estimator $\mathcal{L}^{\infty}$-distance
Wavelet, $k_{0}=0.25$ $0.073$
Wavelet, $k_{0}=0.45$ $0.045$
Wavelet, $k_{0}=0.7$ $0.066$
$\widehat{f}_{s}$, $s=1$ $0.0307$
$\widehat{f}_{s}$, $s=1.25$ $0.0293$
Table 8: The estimated $\mathcal{L}^{\infty}$-distance between the mixture density on $\mathbb{S}^{2}$ and the wavelet estimators and finite-order estimators, for $n=8000$ points.
From this table we can see that our finite-order estimators succeed in achieving a lower $\mathcal{L}^{\infty}$-distance than the wavelet estimators. This is because our estimators are able to simultaneously capture peaks and tails; see the relevant discussion in [1, Section 6].
In summary, our proposed estimators demonstrate that they are flexible enough to estimate multimodal densities with high peaks, a key advantage of adaptive wavelet density estimators, as noted in [1]. Of course, the value of any adaptive estimator is out of any question and is beyond any metrics. The main benefit of our finite-order estimators is their expression, which is more straight-forward to implement and does not require the structural elements behind the wavelet estimators of [1], which are constructed via the theory in [39].
@@ 6.1 Spatially Periodic Data: Honeybee Orientations
In this dataset, bees are labelled as 0, if their orientation is clear, or $1$, if their body is partially obscured within the hive. For consistency in our study, we only look at the $n=691$ honeybee orientations of class 0, which are given in degrees. Converting these to radians makes these observations suitable for a KDE as in Theorem 3.1 and Remark 3.1. We selected $s=1$, and therefore $h=n^{-1/(2s+1)}\approx 0.113$, $N_{s}=14$ and $r=5$. The data and corresponding KDE are plotted in Figure 11. In this frame, the bees exhibit a clear preference for facing approximately towards the angle 0.09-0.09 radians with the least favoured position appearing to be approximately at 2.97 radians.
@@ 6.2 Temporally Periodic Data: Annual Precipitation Patterns
More generally, if a random variable $T$ is periodic on $[a,b]$ with density $f_{T}:[a,b]\to[0,\infty)$ satisfying $f_{T}(a)=f_{T}(b)$, then independent observations $T_{1},\dots,T_{n}\in(a,b]$ may be mapped to circular data on $\mathbb{S}^{1}$ by
$[a,b]$ $\to[-\pi,\pi],$
@@ 6.3 Global Data: Earthquake Locations
With the above conversion in hand, we have observations $X_{1},\dots,X_{1603}\in\mathbb{S}^{2}$. We select $s=0.05$, as the phenomenon is clearly non-uniform and with several zones of high frequency; see Remark 5.1. We construct a KDE $\widehat{f}_{s}$, as in Theorem 3.1 and Remark 3.1, taking $h=n^{-1/(2s+2)}\approx 0.0295$, $N_{s}=92$ and $r=6$. The corresponding KDE as well as the data is plotted in Figure 13.
@@
Country Min. Lat. Max. Lat. Min. Lon. Max. Lon. Freq. KDE Prob.
@@ 6.4 Astronomical Data: Star Directions
Region Min. Lat Max. Lat. Min. Lon. Max. Lon. Freq. KDE Prob.
@@ 7 Conclusion
In our study we provided nonparametric density and probability estimators on the circle $\mathbb{S}^{1}$ and the sphere $\mathbb{S}^{2}$ that are both rate-optimal and computationally efficient. The key idea is to replace the usual infinite expansions, as in [8, 11, 12] underlying kernel density estimators with their finite-order counterparts, determining the order index in accordance with the smoothness of the density. The aforementioned works focused on the theoretical development of density estimation on broad metric spaces and on the further dissemination of the methods used therein. Our present study continues this line of work by focusing on computational aspects, which were not the main focus in [8, 11, 12]. In particular, the need to implement these estimators in practical computations motivated us to investigate theoretically justified truncations of the infinite expansions. The finite-order estimators obtained in Theorem 3.1 preserve the optimality of the previous works and are ready to be used in data analysis.
@@
There are several natural directions for future research. One is the extension of the finite-order methodology to higher-dimensional spheres $\mathbb{S}^{d}$, $d>2$, or other metric spaces associated with Laplacians with discrete spectrum as in [8]. Another natural direction is the use of cross-validation for data-driven selection of the bandwidth, which we plan to investigate in a future work.
@@ Acknowledgements
The authors would like to thank the Editor and the anonymous Referees for their careful reading and constructive comments, which have significantly improved the quality of this manuscript.
The first-named author would like to thank Prof. Morgan Fraser for his helpful comments and suggestions on the astronomical application discussed in Section 6.4.
The second-named author would like to thank Prof. Dr. Katarzyna Bozek an author of [6] and [7], whose dataset and expertise enabled the discussion in zoology in Section 6.1.
@@ 8 Appendix
$\text{MSE}(\widehat{f}_{s})(x)$ $\leq 2\Big(\mathbb{E}_{f}\big[\big(f(x)-\widehat{f}_{n,h}(x)\big)^{2}\big]+\mathbb{E}_{f}\big[\big(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\big)^{2}\big]\Big)$ (8.1)
$=2\Big(\text{MSE}(\widehat{f}_{n,h})(x)+\mathbb{E}_{f}\big[\big(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\big)^{2}\big]\Big).$
@@
$\mathbb{E}_{f}\Big[\big(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\big)^{2}\Big]\leq n^{-2s/(2s+1)}.$
@@ 8.2 Proofs of the results in probability estimation
Then by Jensen’s inequality, for the uniform probability on $A$, and Fubini-Tonelli Theorem we have that
$\mathbb{E}_{f}$ $\big(\widehat{\mathbb{P}}(A)-\mathbb{P}(A)\big)^{2}\leq\mathbb{E}_{f}\Big(\sigma_{\mathbb{S}^{d}}(A)\int_{A}\big|\widehat{f}_{s}(x)-f(x)\big|\frac{d\sigma_{\mathbb{S}^{d}}(x)}{\sigma_{\mathbb{S}^{d}}(A)}\Big)^{2}$
@@
We are concerned with evaluating
$\int_{A[\vartheta_{1},\vartheta_{2}]}\widehat{f}(x)d\sigma_{\mathbb{S}^{1}}(x)=\int_{\vartheta_{1}}^{\vartheta_{2}}\widehat{f}_{s}(\theta)d\theta$
where $\widehat{f}_{s}$ is as in (3.2). By simple integration we confirm the statement:
$\widehat{\mathbb{P}}$ $(A[\vartheta_{1},\vartheta_{2}])=\int_{\vartheta_{1}}^{\vartheta_{2}}\frac{1}{2\pi n}\sum_{j=1}^{n}\Big(1+2\sum_{\ell=1}^{N_{s}}g(h\ell)\cos(\ell(\theta-\theta_{j}))\Big)d\theta$
@@
Therefore for the integral $\int_{\gamma_{1}}^{\gamma_{2}}P_{\ell}^{m}(u)du$, it suffices to compute
$\int_{\gamma_{1}}^{\gamma_{2}}(1+u)^{m/2}(1-u)^{k-m/2}du.$
We consider the more general integral
@@
Re $\big(Y^{m}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{m}(\theta_{j},\phi_{j})}\big)=N_{\ell,m}^{2}P_{\ell}^{m}(\cos\theta)P^{m}_{\ell}(\cos\theta_{j})\text{Re}\Big(e^{im(\phi-\phi_{j})}\Big)$